# Enhancing LLM’s Cognition via Structurization

### v1

@@ Abstract
When reading long-form text, human cognition is complex and structurized. While large language models (LLMs) process input contexts through a causal and sequential perspective, this approach can potentially limit their ability to handle intricate and complex inputs effectively. To enhance LLM’s cognition capability, this paper presents a novel concept of context structurization. Specifically, we transform the plain, unordered contextual sentences into well-ordered and hierarchically structurized elements. By doing so, LLMs can better grasp intricate and extended contexts through precise attention and information-seeking along the organized structures. Extensive evaluations are conducted across various model architectures and sizes (including a series of auto-regressive LLMs as well as BERT-like masking models) on a diverse set of NLP tasks (e.g., context-based question-answering, exhaustive hallucination evaluation, and passage-level dense retrieval). Empirical results show consistent and significant performance gains afforded by a single-round structurization. In particular, we boost the open-sourced LLaMA2-70B model to achieve comparable performance against GPT-3.5-Turbo as the hallucination evaluator. Besides, we show the feasibility of distilling advanced LLMs’ language processing abilities to a smaller yet effective StruXGPT-7B to execute structurization, addressing the practicality of our approach. Code will be made public soon.
@@ 1 Introduction
As illustrated in Fig. 1, the input sequences will be reorganized in a simple but generic three-layer structure: scope, aspects, and descriptions. The scope summarizes the topic and contents of input texts, which is then unfolded into several main aspects with corresponding detailed descriptions. The structurized results can be freely assembled into various natural language forms, depending on the specific downstream tasks. Fig. 2 provides an exemplar of our overall framework: after structurizing and reassembling the vanilla context to highlight its scope and aspects of data imbalance handling techniques, LLMs become able to grasp the target information about the dynamic weighting strategy and generate reliable responses.
@@ 3.1 Efficient Implementation of Structurization
Few-shot commercial LLM. In the initial stage, we use two in-context examples to query the commercial Qwen-Max model to structurize the input corpus, since it shows promising instruction-following and textual analyzing capability with over 200B parameters  [^3] https://rank.opencompass.org.cn/leaderboard-llm-v2. Here is a simplified example of prompting structurization in Fig. 3, and the full template is displayed in Sec. F.1.
@@ 3.2 Effective Utilization of Structurization
After transformation, the natural language input retains its structured knowledge through these systematic cues but is presented in a manner that is inherently more comprehensible for LLMs. This not only facilitates an enhanced understanding and interaction with complex data but also enables the models to leverage their existing natural language capabilities to generate more accurate and contextually relevant responses.
@@ 4 Experiments
Three representative natural language understanding and processing tasks are investigated, including context-based question-answering Sec. 4.1, exhaustive hallucination evaluation Sec. 4.2, and passage-level dense retrieval Sec. 4.3. When instructing tested LLMs to perform target tasks, we merely structurize the vanilla textual inputs, transform the results back into natural languages, and immediately feed the results to LLMs to make responses. The tested LLMs themselves are not fine-tuned on structurized data corpus. We adopt our StruXGPT-7B model to execute structurization in all experiments in this section. The discussion of StruXGPT model is demonstrated in Sec. 4.4.
@@ 4.1 Application on Context-based Question-Answering
Evaluated models. Following [3], we first evaluate three representative large language models on long context comprehension ability: LLaMA2-7B-4k [50], Qwen-7B-8k [2], and ChatGLM3-6B-32k [64]. Then, we extend the evaluation to the larger LLaMA2-13B-4k [50] model. The four LLMs have a relatively similar parameter capacity with different model architectures and window sizes. All the models to examine are pre-trained chat models, and we just employ our StruXGPT to structurize the input contexts to enhance those LLMs’ cognition capabilities. To ensure reproducibility and reduce uncertainty, greedy search is adopted during LLM’s decoding process when generating responses. The accuracy between models’ responses and ground-truth answers is measured by ROUGE-L and F1-score.
@@ 4.2 Application on Exhaustive Hallucination Evaluation
The phenomenon of hallucination in large language models raises doubts about their reliability in generating accurate content [67, 53]. Methods for evaluating hallucinations involve verifying atomic claims against Wikipedia passages [63, 40], yet even advanced LLMs like GPT-3.5-Turbo and GPT-4 cannot always accurately judge the supporting evidence, as LLMs-evaluators often struggle to extract relevant information due to the complexity of passage contexts. We introduce how to improve LLM evaluators’ assessing ability by context structurization below.
@@
Dataset setup. Experiments are conducted on the AttrScore [63] and FactScore [40] datasets. We take the AttrEval-GenSearch test set with approximately 250 examples from AttrScore for evaluation. Each example comprises a statement and its reference passage, and annotators are requested to assign the label of attribution from Attributable (abbreviated as Attr.), Contradictory (abbreviated as Contra.), and Extrapolatory (abbreviated as Extra.). On the other hand, FactScore collected 4,726 atomic claims/statements of people biographies generated by InstructGPT (abbreviated as InstGPT), 5,426 by ChatGPT, and 5,888 by PerplexityAI (abbreviated as PPLAI). For each piece of data in both datasets, we leverage StruXGPT to structurize the reference to identify its main aspects and detailed descriptions. The numerically ordered structure is preserved, as displayed in Fig. A3, since we explicitly ask the evaluator to check the information along the structure for judgment.
Evaluated models. We mainly investigate the open-sourced LLaMA2-7B and LLaMA2-70B models as the LLM-evaluator, and also explore the integration with the close-source GPT-3.5-Turbo-1106 [6] via API access. The main results are presented in Tab. 2. We report the results with Qwen-7B and Qwen-72B models on AttrScore and FactScore in Sec. B.4 for further investigation.
@@ 4.3 Application on Passage-level Dense Retrieval
Retrieval-augmented generation (RAG) has been empirically validated to significantly bolster LLM’s domain knowledge when addressing specialized tasks [27, 59]. During the RAG process, the precision of document retrieval and the in-context learning aptitude of foundational LLMs critically determine the quality of generated outputs. We now investigate the extent to which structurization can facilitate and refine the dense retrieval process for BERT-like masking language models.
@@
Evaluated models. Pre-trained BERT [12], SimLM [54], and coCondenser [15] models are chosen for evaluation, since they achieve the state-of-the-art performance on MS MARCO’s development split. To convert the structurized corpus from our StruXGPT into natural languages, we eliminate the numerical indicators (such as “1.”, “1.1”, etc.) and attach the description statements in the third layer with their aspects. Fig. A3 presents an example. We only structurize the corpus to enhance retrievers’ cognition, and the queries remain unchanged. Following the literature, nDCG@10 is adopted as the evaluation metric, and the results are reported in Tab. 3.
@@
Experimental results. Accordingly, structurization boosts all three retrievers on most of the evaluated datasets, yielding a maximum performance improvement of up to 4.5% on the SciFact subset for SimLM. The results suggest that structurization not only augments decoder-only generative LLMs with explosive parameters (at least 7B), but also benefits encoder-decoder masked language models with constrained parameters (around 110M). It implies that the patterning of linguistic and semantic structurization may be a fundamental mechanism for enhancing language models, transcending distinctions in their architectural design and scale.
4.4 Evaluation of Structurization Approaches
In this section, we assess various structurization methods through exhaustive experiments involving five approaches, including prompting Qwen-max with few-shot exemplars (serves as our teacher model), few-shot results from Qwen-7B and LLaMA2-7B pre-trained chat models, and our fine-tuned Qwen-7B and LLaMA2-7B (student) models. As introduced in Sec. 3.1, we use 200 validation cases and have the five models generate 1,000 structurized outputs for analysis.
@@ 4.4 Evaluation of Structurization Approaches
The commercial Qwen-max shows a promising instruction-following and in-context learning ability, generally scoring 4.5 at the three dimensions. However, both the pre-trained 7B-parameter chat models from Qwen and LLaMA2 immediately decline the scores to below 4.0, as they struggle to understand the instruction to build the three-layer structure and tend to hallucinate responses due to the limited model capacity. More structurization examples can be found in Sec. F.2. Notably, the fine-tuned StruXGPT-7B-Q(wen) and StruXGPT-7B-L(LaMA2) both obtain a 4.35 - 4.45 score on average. They inherit 97% of structurization capability from the Qwen-max teacher model, evidencing the effectiveness of training a specialized 7B-parameter model for structurization to aid in efficiency and privacy.
@@
Evaluation with semantic embedding. (EmbEval) At last, we explore the structurization quality evaluation in the semantic embedding perspective, as a supplementary to lexical evaluation. Following Zhang et al. [65], we calculate the semantic similarity between original and structurized contents with the embedding similarities, and the results in Tab. 4 shows consistent measure against HumanEval and AppEval with a much lower cost. Hence, BERTScore [65] can be further leveraged as an effective and efficient quality-assessment tool to filter training data for StruXGPT to achieve better structurization. We view this as our future work.
@@ 5 Conclusion and Discussion
This paper presents a novel concept of context structurization to enhance LLM’s cognition capability. Extensive evaluations are performed on various representative NLP tasks, and the results reveal structurization’s consistent effectiveness across the base language model’s architectural designs and capacity scales. In addition, we also demonstrate the feasibility of distilling the context structurization ability from giant commercial LLMs into a responsive and private StruXGPT-7B model, addressing the practicality of our approach. The limitation and future work are discussed in Appendix C. We hope this paper can bring new insights into how to build a more powerful and reliable language model with human cognition and intelligence.
@@ Appendix A Implementation Details of Structurization
A.1 Training Details
The data to train our StruXGPT is carefully curated from two main aspects. First, to ensure diversity, we randomly sample 15,000 passages from Wikipedia across various domains, whose token length ranges from 1K to 3K. Then, we also supply round 15,000 statement passages from the domain-specific model- and human-generated content, including the filtered CAMEL-AI [29], FiQA [38], and MedQuad [4] datasets. After calling the Qwen-max API, we collect the structurized results and eliminate the failed cases (usually due to internet inaccessibility and unexpected output format), resulting in 22,547 pieces of training data in total to derive our StruXGPT, and 200 test samples for evaluation.
StruXGPT is built upon LLaMA2- or Qwen-7B-chat models, which had been aligned to humans with promising instruction-following capabilities with their knowledge prior. The 22,547 pieces of input-output pairs are utilized to distill the structurization ability to StruXGPT via supervised fine-tuning (SFT). Specifically, StruXGPT is trained with a constant learning rate of $5\times 10^{-6}$ for LLaMA and $1\times 10^{-5}$ for Qwen for 1 epoch. The batch size is 128, and other hyper-parameters follow the default settings from Touvron et al. [50] and Bai et al. [2]. The training is resource-friendly, which is done on 8 NVIDIA V100 (16G) GPUs for 3.5 hours.

### v2

@@ Abstract
When reading long-form text, human cognition is complex and structurized. While large language models (LLMs) process input contexts through a causal and sequential perspective, this approach can potentially limit their ability to handle intricate and complex inputs effectively. To enhance LLM’s cognition capability, this paper presents a novel concept of context structurization. Specifically, we transform the plain, unordered contextual sentences into well-ordered and hierarchically structurized elements. By doing so, LLMs can better grasp intricate and extended contexts through precise attention and information-seeking along the organized structures. Extensive evaluations are conducted across various model architectures and sizes (including a series of auto-regressive LLMs as well as BERT-like masking models) on a diverse set of NLP tasks (e.g., context-based question-answering, exhaustive hallucination evaluation, and passage-level dense retrieval). Empirical results show consistent and significant performance gains afforded by a single-round structurization. In particular, we boost the open-sourced LLaMA2-70B model to achieve comparable performance against GPT-3.5-Turbo as the hallucination evaluator. Besides, we show the feasibility of distilling advanced LLMs’ language processing abilities to a smaller yet effective StruXGPT-7B to execute structurization, addressing the practicality of our approach. Code is available at https://github.com/alibaba/struxgpt.
@@ 1 Introduction
As Fig. 1 illustrates, input sequences are reorganized in a simple but generic three-layer structure: scope, aspects, and descriptions. The scope summarizes the topic and contents, unfolding into several main aspects with corresponding detailed descriptions. The structurized results can be freely assembled into various natural language forms, depending on the specific downstream tasks. Fig. 2 provides an exemplar of our overall framework: after structurizing and reassembling the vanilla context to highlight its scope and aspects of data imbalance handling techniques, LLMs become able to grasp the target information about the dynamic weighting strategy and generate reliable responses.
@@ 3.1 Efficient Implementation of Structurization
Few-shot commercial LLM. In the initial stage, we use two in-context examples to query the commercial Qwen-Max model to structurize the input corpus, since it shows promising instruction-following and textual analyzing capability with over 200B parameters  [^3] https://rank.opencompass.org.cn/leaderboard-llm-v2. Here is a simplified example of prompting structurization in Fig. 3, and the full template is displayed in Sec. E.1.
@@ 3.2 Effective Utilization of Structurization
After transformation, the linguistic input retains its structured knowledge through systematic cues but is presented in a comprehensible manner for LLMs. This not only facilitates an enhanced understanding and interaction with complex data but also enables the models to leverage their existing natural language capabilities to generate more accurate and contextually relevant responses.
@@ 4 Experiments
Three representative natural language understanding and processing tasks are investigated, including context-based question-answering Sec. 4.1, exhaustive hallucination evaluation Sec. 4.2, and passage-level dense retrieval Sec. 4.3. When instructing tested LLMs to perform target tasks, we merely structurize the vanilla textual inputs, transform the results back into natural languages, and immediately feed the results to LLMs to make responses. The tested LLMs themselves are not fine-tuned on structurized data corpus. We adopt our StruXGPT-7B model to execute structurization in all experiments in this section. The evaluation and ablation of StruXGPT model are demonstrated in Sec. 4.4, Sec. 4.5, and Sec. 4.6, and more comparison with related augmentation-based methods can be found in Sec. B.3 and Sec. B.5.
@@ 4.1 Application on Context-based Question-Answering
Evaluated models. Following [3], we evaluate three representative large language models on long context comprehension ability: LLaMA2-7B-4k [53], Qwen-7B-8k [2], and ChatGLM3-6B-32k [67], and extend the larger LLaMA2-13B-4k [53] model. The four LLMs have a relatively similar parameter capacity with different model architectures and window sizes. All the models to examine are pre-trained chat models, and we just employ our StruXGPT to structurize the input contexts to enhance those LLMs’ cognition capabilities. To ensure reproducibility and reduce uncertainty, greedy search is adopted during LLM’s decoding process when generating responses. The accuracy between models’ responses and ground-truth answers is measured by ROUGE-L and F1-score.
@@ 4.2 Application on Exhaustive Hallucination Evaluation
Hallucination has raised wide attention in the community [70, 56]. In general, evaluating hallucinations involves verifying atomic claims against supportive materials (e.g., Wikipedia passages [66, 42]), yet even advanced GPT-3.5-Turbo and GPT-4 cannot always accurately make the judge, as LLMs-evaluators often struggle to extract relevant information due to the complexity of passage contexts. We introduce how to improve LLM evaluators’ assessing ability by context structurization below.
@@
Dataset setup. AttrScore [66] and FactScore [42] datasets are adopted for evaluation. We take the AttrEval-GenSearch test set with 245 examples from AttrScore, where each example comprises a statement and its reference passage, and is annotated by Attributable (abbreviated as Attr.), Contradictory (abbreviated as Contra.), and Extrapolatory (abbreviated as Extra.). FactScore collected 4,726 atomic claims/statements of people biographies generated by InstructGPT (abbreviated as InstGPT), 5,426 by ChatGPT, and 5,888 by PerplexityAI (abbreviated as PPLAI). For each input sample, we leverage StruXGPT to structurize the reference to identify its main aspects and detailed descriptions. The numerically ordered structure is preserved, as displayed in Fig. A3, since we explicitly ask the evaluator to check the information along the structure for judgment.
Evaluated models. We mainly investigate the open-sourced LLaMA2-7B and LLaMA2-70B models as the LLM-evaluator, and also explore the integration with the close-source GPT-3.5-Turbo-1106 [6] via API access. The main results are presented in Tab. 2. We also report the results with Qwen models on AttrScore and FactScore and the incorporation to more powerful GPT-4 in Sec. B.4.
@@ 4.3 Application on Passage-level Dense Retrieval
Retrieval-augmented generation (RAG) has been empirically validated to significantly bolster LLM’s domain knowledge [29, 62], where precise document retrieval plays a vital role. We now investigate how structurization can facilitate dense retrieval for BERT-like masking language models.
@@
Evaluated models. BERT [13], SimLM [57], and coCondenser [16] are chosen for evaluation, since they achieve state-of-the-art performance on MS MARCO’s development split. To convert the structurized passages from our StruXGPT into natural languages, we eliminate the numerical indicators (such as “1.”, “1.1”, etc.) and attach the description statements in the third layer with their aspects. Fig. A3 presents an example. We only structurize passages to enhance retrievers’ cognition, and the queries remain unchanged. Following the literature, nDCG@10 results are reported in Tab. 3.
@@
Experimental results. Structurization boosts all three retrievers on most subsets, yielding a maximum performance improvement of 4.5% on SciFact for SimLM. The results suggest that structurization not only augments decoder-only generative LLMs with explosive parameters (at least 7B), but also benefits encoder-decoder masked language models with constrained parameters (around 110M). It implies that the patterning of linguistic and semantic structurization may be a fundamental mechanism for enhancing language models, transcending distinctions in their architectural design and scale.
4.4 Evaluation of the Structurization Approach Itself
In this section, we assess various structurization methods through exhaustive experiments on five approaches, including prompting Qwen-max with few-shot exemplars (serves as our teacher model), few-shot Qwen-7B and LLaMA2-7B pre-trained chat models, and our fine-tuned Qwen-7B and LLaMA2-7B (student) models. As introduced in Sec. 3.1, we use 200 validation cases and have the five models generate 1,000 structurized outputs for analysis.
@@ 4.4 Evaluation of the Structurization Approach Itself
The commercial Qwen-max shows a promising instruction-following and in-context learning ability, generally scoring 4.5 at the three dimensions. However, the pre-trained 7B models from Qwen and LLaMA2 immediately decline the scores to below 4.0, as they struggle to understand the instruction to build the three-layer structure and tend to hallucinate responses due to the limited model capacity. More structurization examples can be found in Sec. E.2. Notably, the fine-tuned StruXGPT-7B-Q(wen) and StruXGPT-7B-L(LaMA2) both obtain a 4.35 - 4.45 score on average. They inherit 97% of structurization capability from the Qwen-max teacher model, evidencing the effectiveness of training a specialized 7B-parameter model for structurization to aid in efficiency and privacy.
@@
Evaluation with semantic embedding (SemEval). At last, we explore the structurization quality evaluation in the semantic embedding perspective, as a supplementary to lexical evaluation. Following Zhang et al. [68], we calculate the semantic similarity between original and structurized contents with the embedding similarities, and the results in Tab. 4 show consistent measure against HumanEval and AppEval with a much lower cost. Hence, BERTScore [68] can be further leveraged as an effective and efficient quality-assessment tool for training-data filtering and structurization quality evaluation to derive a better StruXGPT model. Sec. 4.6 presents some preliminary investigations.
@@
4.5 Ablation Studies on StruXGPT’s Establishment
This section studies two major factors of training a StruXGPT model: data quality and model capacity.
Using two few-shot examples is sufficient to collect high-quality training data. In this work, we choose 2 in-context examples to prompt commercial LLMs (as a teacher) to generate data pairs of raw/structurized texts to train our StruXGPT-7B model (as a student). We think it is enough for teacher models to understand the structurization process and generate valid training samples, as the 2 examples respectively describe the 2 most common types of real-world text (i.e., with/without existing indicators like “1.”, “2.”, etc), which is displayed in Sec. E.2.
To further verify it, we investigate the number of in-context examples with two evaluation protocols (as in Tab. 4): AppEval (an improvement on Qasper subset with context structurization) and BERTScore (semantic similarity with raw and structurized texts in the validation set). We also report the error rate when parsing structurization results from the teacher model’s outputs (denoted as “FormatError”).
According to Sec. 4.5, 1-shot is apparently insufficient to illustrate structurization, while 2- and 3-shot achieve comparable structurization quality under AppEval and BERTScore. Notably, 3-shot receives a 2% lower FormatError than 2-shot, in trade for the increased inference cost (because of increased few-shot samples). We argue that the 2% gap (around 400 samples) does not make a difference for the final StruXGPT training, which can be verified in Sec. A.2. Therefore, we recommend users to apply 3- or even more shots when prompting teacher LLMs if available, otherwise 2-shot is also a good choice to balance the inference cost and structurization quality.
Table 5: Number of few-shot examples.
nShot AppEval BERTScore FormatError
1-shot +1.8 0.282 25.4%
2-shot +3.2 0.308 7.4%
3-shot +3.3 0.302 5.5%
Table 6: Parameter capacity of StruXGPT.
StruXGPT AppEval BERTScore FormatError
Qwen-1.8B +2.7 0.299 5.0%
Qwen-7B +3.6 0.313 0.0%
Qwen-14B +3.8 0.323 0.0%
StruXGPT-7B balances parameter capacity and structurization quality. As Qwen [2] provides a series of models varying sizes, in Sec. 4.5, we implemented StruXGPT on Qwen-1.8B/7B/14B respectively to investigate the relationship between model capacity and structurization quality.
Compared with the 7B model capacity, the smaller 1.8B model, despite its positive enhancement on downstream applications, shows slight inferiority in both AppEval (+2.7 v.s. +3.6) and BERTScore (0.299 v.s. 0.313), and presents 5% error rate when parsing structurization results. On the other hand, the 14B model brings further improvement to BERTScore, (the structurization content is relatively more faithful to original texts), but the boost on AppEval is insignificant. Hence, the 7B model is a good trade-off between model capacity (training/inference efficiency) and structurization quality.
4.6 Utilization of Structurization Quality Assessment
As the structurization quality simultaneously influences StruXGPT’s training performance and application improvements, this section investigates the quality assessment tool BERTScore (as discussed in Sec. 4.2) on StruXGPT’s training and inference stages, respectively.
On one hand, as the training data quality determines StruXGPT’s upper bound, we statistic the BERTScore of the 22K training entries, and around 94.45% raw/structurized text pairs present positive scores (normalized by the baseline score of 0.83, and a positive BERTScore presents a benign similarity), demonstrating the high quality of our training data. Consequently, we eliminated around 5% of data with negative scores and trained another StruXGPT model, and the results in Sec. 4.6 indicate this part of data does not affect the final performance. According to Sec. A.2, data quantity may play a vital role in further improving our StruXGPT.
Table 7: Training data filtering.
Training Data AppEval BERTScore
vanilla +3.6 0.313
filtered +3.4 0.316
Table 8: Inference results filtering.
Context Declined Ratio Overall Enhance
vanilla 3.5% +3.6
filtered 3.0% +3.7
On the other hand, since poor structurization results can lead to suboptimal application performance, we statistic the enhancement variance in the Qasper subset from LongBench [3] in Sec. 4.6 for structurized context inputs. Our method merely causes degradation on 3.5% of samples (with relatively lower structurization quality), but ultimately receives a +3.6 improvement overall test samples. Furthermore, we filter out the structurization results with a low BERTScore (e.g., $<$0.05) and take back the original context as input. In this way, the degradation can be alleviated (from 3.5% to 3.0%), further improving the final enhancement to +3.7. The lower bound of our StruXGPT can thus be ensured.
@@ 5 Conclusion and Discussion
This paper presents a novel concept of context structurization to enhance LLM’s cognition capability. Extensive evaluations of various representative NLP tasks reveal the consistent enhancement across language model’s architectural designs and capacity scales. We demonstrate the feasibility of distilling the structurization ability from giant commercial LLMs into a responsive and private StruXGPT-7B model, addressing the practicality problem. The limitation and future work are discussed in Appendix C. We hope this paper can bring new insights into how to build a more powerful and reliable language model with human cognition and intelligence.
Acknowledgments and Disclosure of Funding
This work was supported in part by the Fundamental Research Funds for the Central Universities, in part by Alibaba Cloud through the Research Intern Program, and in part by Zhejiang Provincial Natural Science Foundation of China under Grant No. LDT23F01013F01.
@@ Appendix A Implementation Details of Structurization
A.1 Implementation Details
The data to train our StruXGPT is carefully curated from two main aspects. First, to ensure diversity, we randomly sample 15,000 passages from Wikipedia across various domains, whose token length ranges from 1K to 3K. Then, we also supply round 15,000 statement passages from the domain-specific model- and human-generated content, including the filtered CAMEL-AI [31], FiQA [40], and MedQuad [4] datasets. With Qwen-max API, we collect the structurized results and eliminate the failed cases (usually due to internet inaccessibility and unexpected output format), resulting in 22,547 pieces of training data in total to derive our StruXGPT, and 200 test samples for evaluation.
StruXGPT is built upon LLaMA2- or Qwen-7B-chat models, which had been aligned to humans with promising instruction-following capabilities with their prior knowledge. The 22,547 pieces of input-output pairs are utilized to distill the structurization ability to StruXGPT via supervised fine-tuning (SFT). Specifically, StruXGPT is trained with a constant learning rate of $5\times 10^{-6}$ for LLaMA and $1\times 10^{-5}$ for Qwen for 1 epoch. The batch size is 128, and other hyper-parameters follow the default settings from Touvron et al. [53] and Bai et al. [2]. The training is resource-friendly, which can be done on 8 NVIDIA V100 (16G) GPUs for 3.5 hours.
For all the inference experiments, we leverage 1-2 NVIDIA A100-80G GPUs for model deployment.